OpenStudy (anonymous):

You can walk 2.5 miles per hour. How many hours will it take you to walk 12.5 miles?  

OpenStudy (anonymous):

5 hours

OpenStudy (anonymous):

i personly just did this mentaly and i think its 6 hrs

OpenStudy (anonymous):

4 * 2.5 = 10 so 5 * 2.5 = 12.5

OpenStudy (anonymous):

i just did 6\[\times\] 2 = 12

OpenStudy (anonymous):

grr i meant 6x2=12 so yah

OpenStudy (anonymous):

6*2.5 = 15 (6*2) + (6*.5) = 12 + 3 = 15

OpenStudy (anonymous):

k im gonna shut up now